The following are the prices in Rs. of different brands of television sets. Find the range of



prices and its coefficient.



Rs. 17850, 16990, 17500, 19850, 16650, 19300

1
Expert's answer
2021-12-21T01:55:03-0500

The following are the prices in Rs. of different brands of television sets:

(17850, 16990, 17500, 19850, 16650, 19300)

Range = Maximum Price – Minimum Price

Range = 19850 - 16650

Range = 3200

So, the range of prices is Rs. 3200.


"Coefficient=\frac{Maximum-Minimum}{Maximum+Minimum}"


"Coefficient=\frac{19850 - 16650 }{19850 + 16650 }"


"Coefficient=0.088"